OpenStudy (ttop0816):

What is the magnitude of the position vector whose terminal point is (-5, 3)? 2 sqrt 17 4 5 6

OpenStudy (sidsiddhartha):

magnitude of a vector is the modulus value like for a vector--(a,b) its magnitude will be \[mag(a,b)=\sqrt{a^2+b^2}\]

OpenStudy (sidsiddhartha):

can u do it now?

OpenStudy (ttop0816):

is it 6?

OpenStudy (sidsiddhartha):

nope (-5,3) can be represented as\[V=-5 \hat i+3 \hat j\] so its magnitude will be \[|v|=\sqrt{(-5)^2+3^2}=\sqrt{25+9}=\sqrt{34}\]

OpenStudy (sidsiddhartha):

got that?

OpenStudy (phi):

Based on the answer choices, this question has a typo. If they said the vector had components (-4,3) it would work out. or if then included sqr(34) as a choice... but as it stands, you should mention to your teacher there is a problem with this question.

OpenStudy (ttop0816):

yes!

OpenStudy (ttop0816):

then it would be 5 because square root 34= 5.830?

OpenStudy (sidsiddhartha):

5.830 is closer to 6 so it should be 6 if sqrt{34} is'nt any option